### v1

# 2D-OOB: Attributing Data Contribution through Joint Valuation Framework

Abstract
Data valuation has emerged as a powerful framework to quantify the contribution of each datum to the training of a particular machine learning model. However, it is crucial to recognize that the quality of various cells within a single data point can vary greatly in practice. For example, even in the case of an abnormal data point, not all cells are necessarily noisy. The single scalar valuation assigned by existing methods blurs the distinction between noisy and clean cells of a data point, thereby compromising the interpretability of the valuation. In this paper, we propose 2D-OOB, an out-of-bag estimation framework for jointly determining helpful (or detrimental) samples, as well as the particular cells that drive them. Our comprehensive experiments demonstrate that 2D-OOB achieves state-of-the-art performance across multiple use cases, while being exponentially faster. 2D-OOB excels in detecting and rectifying fine-grained outliers at the cell level, as well as localizing backdoor triggers in data poisoning attacks.
1 Introduction
From customer behavior prediction and medical image analysis to autonomous driving and policy making, machine learning (ML) systems process ever increasing amounts of data. In such data-rich regimes, a fraction of the samples is often noisy, incorrect annotations are likely to occur, and uniform data quality standards become difficult to enforce. To address these challenges, data valuation emerges as a research field receiving increasing attention, focusing on properly assessing the contribution of each datum to ML training [12]. These methods have proven useful in identifying low-quality samples that can be detrimental to model performance, as well as selecting subsets of data that are representative of enhanced model performance [23, 48, 27]. Furthermore, they are widely applicable in data marketplace for fair revenue allocation and incentive design [51, 45, 40].
Nevertheless, existing data valuation methods assign a scalar score to each datum, thereby failing to account for the varied roles of individual cells. This leaves the valuation rationale unclear and can be unsatisfactory and sub-optimal in various practical scenarios. Firstly, whenever a score is assigned to a data point by a particular data valuation method, it is crucial to understand the underlying justifications to ensure transparency and reliability, especially in high-stakes decision making [39]. Secondly, it is important to recognize the fact that even if a data point is of low quality, it is rarely the case that all the cells within this data point are noisy [37, 26, 43]. The absence of detailed insights into how individual cells contribute to ML training inevitably leads to discarding entire data points. This can result in substantial data waste, particularly when only a few cells are noisy and data acquisition is expensive. Finally, in data markets, different cells within a data point may originate from different data sellers [3, 10]. Consequently, a singular valuation for the entire point fails to offer equitable compensation to all contributing parties.
Figure 1: Comparison of data valuation and joint valuation. (a) Data valuation evaluates the quality of individual data points, whereas (b) joint valuation evaluates the quality of individual cells. Both panels illustrate the same hypothetical dataset, and the darker colors overlaid represent the higher quality or importance. Joint valuation provides a finer level of attributions than data valuation and aims to describe how features affect data values. As panel (b) illustrates, the joint valuation framework can identify outlier cells highlighted with blue boxes (i.e., $-1$ in "Income" and $100$ in "Education") and provide quantitative interpretations of data values.
Our contributions
In this paper, we propose 2D-OOB, a powerful and efficient joint valuation framework that can attribute a data point’s value to its individual features. 2D-OOB quantifies the importance of each cell in a dataset, as illustrated in Figure 1, providing interpretable insights into which cells are associated with influential data points. Our method is computationally efficient as well as theoretically supported by its connections with Data-OOB [23]. Moreover, our extensive empirical experiments demonstrate the practical effectiveness of 2D-OOB in various use cases. 2D-OOB accurately identifies cell outliers and pinpoints which cells to fix to improve model performance. 2D-OOB enables inspection of data poisoning attacks by precisely localizing the backdoor trigger, an artifact inserted into a training sample to induce malicious model behavior [13, 5]. 2D-OOB is on average $200$ times faster than state-of-the-art methods across all datasets examined.
2 Preliminaries
Notations
Throughout this paper, we focus on supervised learning settings. For $d\in\mathbb{N}$, we denote an input space and an output space by $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}\subseteq\mathbb{R}$, respectively. We denote a training dataset with $n$ data points by $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ where $(x_{i},y_{i})$ is the $i$-th pair of the input covariates $x_{i}\in\mathcal{X}$ and its output label $y_{i}\in\mathcal{Y}$. For an event $A$, an indicator function $\mathds{1}(A)$ is $1$ if $A$ is true, otherwise 00. For $j\in\mathbb{N}$, we set $[j]:=\{1,\dots,j\}$. For a set $S$, we denote its power set by $2^{S}$ and its cardinality by $|S|$.
DataShapley
The primary goal of data valuation is to quantify the contribution of individual data points to a model’s performance. Leveraging the Shapley value in cooperative game theory [38], DataShapley [12] measures the average change in a utility function $U:2^{\mathcal{D}}\rightarrow\mathbb{R}$ when a data point is removed. For $i\in[n]$, DataShapley of $i$-th datum is defined as follows.
$\phi_{i}^{\mathrm{Shap}}:=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\binom{n-1}{k-1}}\sum_{S\subset\mathcal{D}_{k}^{(i)}}[U(S\cup\{(x_{i},y_{i})\})-U(S)]$ (1)
where $\mathcal{D}_{k}^{(i)}:=\{S\subseteq\mathcal{D}|(x_{i},y_{i})\notin S,|S|=k-1\}$. DataShapley $\phi_{i}^{\mathrm{Shap}}$ in (1) considers every set $S\in\mathcal{D}_{k}^{(i)}$ and computes the average difference in utility $U(S\cup\{(x_{i},y_{i})\})-U(S)$. It characterizes the impact of a data point, but its computation requires evaluating $U$ for all possible subsets of $\mathcal{D}$, rendering precise calculations infeasible. Many efficient computation algorithms have been studied [15, 25, 50], and in these studies, Shapley-based methods have demonstrated better effectiveness in detecting low-quality samples than standard attribution approaches, such as leave-one-out and influence function methods [20, 9].
Data-OOB
As an alternative efficient data valuation method, Kwon and Zou [23] propose Data-OOB, which leverages a bagging model and measures the similarity between a nominal label and weak learners’ predictions. To be more specific, we suppose a bagging model consists of $B$ weak learners, where for $b\in[B]$, the $b$-th weak learner $\hat{h}_{b}$ is given as a minimizer of the weighted empirical risk,
$\hat{h}_{b}:=\mathrm{argmin}_{h}\sum_{i=1}^{n}w_{bi}\ell(y_{i},h(x_{i})),$
where $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a loss function and $w_{bi}\in\mathbb{N}$ is the number of times the $i$-th datum $(x_{i},y_{i})$ is selected by the $b$-th bootstrap dataset. Let $\mathbf{w}_{b}$ be a weight vector $\mathbf{w}_{b}:=(w_{b1},\dots,w_{bn})$ for all $b\in[B]$. For $i\in[n]$ and $\{(\mathbf{w}_{b},\hat{h}_{b})\}_{b=1}^{B}$, Data-OOB of the $i$-th datum is defined as follows.
$\phi_{i}^{\mathrm{OOB}}:=\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)T(y_{i},\hat{h}_{b}(x_{i}))}{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)},$ (2)
where $T(y_{i},\hat{h}_{b}(x_{i}))$ is a score function evaluated at $(x_{i},y_{i})$. We assume that the higher $T$, the better the prediction. In classification settings, a common choice for $T$ is $\mathds{1}(y_{i}=\hat{h}_{b}(x_{i}))$, and in this case, Data-OOB $\phi_{i}^{\mathrm{OOB}}$ measures the average similarity between a nominal label $y_{i}$ and weak learners’ predictions $\hat{h}_{b}(x_{i})$ when a datum $(x_{i},y_{i})$ is not sampled in a bootstrap dataset. It intuitively captures the quality of a data point. For instance, when $(x_{i},y_{i})$ is a mislabeled sample or an outlier, the label $y_{i}$ is likely to differ from $\hat{h}_{b}(x_{i})$, resulting in $\phi_{i}^{\mathrm{OOB}}$ being close to zero.
It is noteworthy that Data-OOB in (2) can be computed by training a single bagging model, making it computationally efficient. Kwon and Zou [23] show that Data-OOB can easily scale to millions of data points, but for DataShapley this is often very impractical. In addition, Data-OOB is often comparable to or even more effective than DataShapley in detecting mislabeled data points and selecting helpful data points [23, 16].
3 Attributing Data Contribution through Joint Valuation Framework
Data valuation quantifies desiderata of data points, however, it does not describe what features contribute and how much to those specific data values. For instance, in anomaly detection tasks, data valuation methods can be deployed to detect anomalous data points, but they do not explain why they are abnormal, which is not generally desirable in practice. To address this challenge, we consider a joint valuation framework and assess a cell score for each feature of a data point. Here, a cell score is designed to quantify how a feature affects the value of an individual data point, attributing a data value to features.
To the best of the author’s knowledge, Liu et al. [29] first consider a concept of the joint valuation in literature and introduce 2D-Shapley to quantitatively interpret DataShapley. To this end, we denote a 2D utility function by $u:[n]\times[d]\to\mathbb{R}$, which takes as input a subset of data points $S\subseteq[n]$ and a subset of features $F\subseteq[d]$, and measure the utility of a fragment of the given dataset consisting of cells $\{(i,j)\}_{i\in S,j\in F}$, where a tuple $(i,j)$ denotes a cell at the $i$-th datum and the $j$-th column. Then, 2D-Shapley is defined as
$\psi_{ij}^{\mathrm{2D-Shap}}:=\frac{1}{nd}\sum_{k=1}^{n}\sum_{l=1}^{d}\frac{1}{\binom{n-1}{k-1}\binom{d-1}{l-1}}\sum_{(S,F)\subset\mathcal{D}_{k,l}^{(i,j)}}M_{u}^{i,j}(S,F)$ (3)
where $\mathcal{D}_{k,l}^{(i,j)}:=\{(S,F)|S\subseteq[n]\backslash\{i\},F\subseteq[d]\backslash\{j\},|S|=k-1,|F|=l-1\}$ and
$M_{u}^{i,j}(S,F)=u(S\cup\{i\},F\cup\{j\})+u(S,F)-u(S\cup\{i\},F)-u(S,F\cup\{j\}).$
The function $M_{u}^{i,j}$ allows us to quantify how much removing a specific cell at $(i,j)$ from a given set $(S\cup\{i\},F\cup\{j\})$ affects the overall utility, and 2D-Shapley $\psi_{ij}^{\mathrm{2D-Shap}}$ evaluates the average $M_{u}^{i,j}$ across all possible data fragments $(S,F)\subset\mathcal{D}_{k,l}^{(i,j)}$.
Similar to DataShapley, the permutation of all rows and columns required for exact 2D-Shapley calculations presents significant computational challenges. To address this, Liu et al. [29] develop 2D-KNN, which utilizes $k$-nearest-neighbors models as surrogates to approximate 2D-Shapley values. However, the approximation methods can compromise the accuracy of valuations [23, 16]. Additionally, 2D-KNN still faces challenges scaling to large-scale datasets and high-dimensional settings.
We propose 2D-OOB, an efficient and model-agnostic joint valuation framework that leverages out-of-bag estimation to attribute data contribution. We further illustrate how 2D-OOB is connected to Data-OOB, thereby facilitating sample-wise interpretation for data valuation in Section 3.2.
3.1 2D-OOB: an efficient joint valuation framework
Our idea builds upon the subset bagging model [14], which is well recognized as an earlier version of Breiman’s random forest model [4]. A key distinction from a standard bagging model is that a weak learner in a subset bagging model is trained on a randomly selected subset of features. For $b\in[B]$, we denote the $b$-th random feature subset by $S_{b}\subseteq[d]$. Then, the $b$-th weak learner of a subset bagging model is given as follows.
$\hat{f}_{b}:=\mathrm{argmin}_{f}\sum_{i=1}^{n}w_{bi}\ell(y_{i},f(x_{i,S_{b}})),$
where $x_{i,S_{b}}$ is a subvector of $x_{i}$ that only takes elements in a subset $S_{b}$. This difference enables us to assess the impact of which features are more influential: if $S_{b}$ includes a helpful (or detrimental) feature, we can expect the out-of-bag prediction $\hat{f}(x_{i,S_{b}})$ to be good (or poor). We formalize this intuition and propose 2D-OOB. For $i\in[n]$, $j\in[d]$ and $\{(w_{b},S_{b},\hat{f}_{b})\}_{b=1}^{B}$, the 2D-OOB for the $j$-th cell of the $i$-th data point is defined as follows,
$\psi_{ij}^{\mathrm{2D-OOB}}:=\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0,j\in S_{b})T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))}{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0,j\in S_{b})},$ (4)
where $T:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a utility function that scores the performance of the weak learner $\hat{f}_{b}(x_{i,S_{b}})$ on the $i$-th datum $(x_{i},y_{i})$. Specifically, for binary or multi-class classification problems, we can adopt $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}_{b}(x_{i,S_{b}}))$. In this case, 2D-OOB measures the average accuracy score of out-of-bag predictions (specifically, when the $i$-th data point is out-of-bag) if a cell $j$ is used in training $\hat{f}_{b}$. For regression problems, we can use the negative squared error loss function, defined as $T(y_{i},\hat{f}_{b}(x_{i,S_{b}}))=-(y_{i}-\hat{f}_{b}(x_{i,S_{b}}))^{2}$. In practice, $\mathcal{X}$ could also be incorporated into $T$ to suit the specific use case.
While Data-OOB in (2) aims to assess the impact of the $i$-th datum, 2D-OOB in (4) provides interpretable insights by evaluating the data point with various combinations of features, revealing which cells are influential to model performance. Leveraging subset bagging scheme, 2D-OOB requires a single training of the bagging model, and thus it is computational efficiency.
3.2 Connection to Data-OOB
We now present interpretable expressions of how 2D-OOB connects to Data-OOB in the following proposition. To begin with, we denote a set of subsets of $[d]$ by $\mathcal{S}:=\{S\subseteq[d]\}$. With $\{(\mathbf{w}_{b},\hat{f}_{b})\}_{b=1}^{B}$, we define the $i$-th Data-OOB when a particular subset $S$ is used as follows and denote it by $\phi_{i}^{\mathrm{OOB}}(S)$.
$\phi_{i}^{\mathrm{OOB}}(S):=\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)T(y_{i},\hat{f}_{b}(x_{i,S}))}{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)}.$
Proposition 3.1.
For all $i\in[n]$ and $j\in[d]$, $\psi_{ij}^{\mathrm{2D-OOB}}$ can be expressed as follows.
$\psi_{ij}^{\mathrm{2D-OOB}}$ $=\mathbb{E}_{\hat{F}_{S}}[\phi_{i}^{\mathrm{OOB}}(S)\mid j\in S],$
where $\hat{F_{S}}$ is an empirical distribution with respect to $S$ induced by the sampling process.
A proof is given in the Appendix C. Proposition 3.1 shows that 2D-OOB $\psi_{ij}^{\mathrm{2D-OOB}}$ can be expressed as a conditional empirical expectation of Data-OOB provided that the $j$-th feature is used in Data-OOB computation. It provides intuitive interpretations: for a fixed $i$ and $j\neq k$, $\psi_{ij}^{\mathrm{2D-OOB}}>\psi_{ik}^{\mathrm{2D-OOB}}$ implies that the cell $x_{ij}$ is more helpful to achieve the high OOB score, which serves as an indicator of model performance, than the cell $x_{ik}$.
Figure 2: Cell-level outlier detection rate curves for 2D-OOB, 2D-KNN, and Random. The x-axis represents the percentage of inspected cells. The y-axis represents the detection rate, defined as the ratio of the number of detected outlier cells to the total number of outlier cells present in a dataset. The error bars show a $95\%$ confidence interval based on $30$ independent experiments. We examine the cells in ascending order, starting from those with the lowest values, and thus a curve closer to the left-top corner indicates better performance. 2D-OOB efficiently detects the majority of outlier cells by examining only a small fraction of the total cells, while 2D-KNN and Random require scanning nearly all the cells.
4 Experiments
In this section, we empirically show the effectiveness of 2D-OOB across multiple use cases of the joint valuation: cell-level outlier detection, cell fixation, and backdoor trigger detection. As a summary, 2D-OOB can precisely identify anomalous cells that should be prioritized for examination and subsequent fixation to improve model performance. In the context of backdoor trigger detection, 2D-OOB demonstrates its efficacy by accurately identifying different types of triggers within poisoned data, showcasing its proficiency in detecting non-random, targeted anomalies. Our method also exhibits high computational efficiency through run-time comparison.
Throughout all of our experiments, 2D-OOB uses a subset bagging model with $B=1000$ decision trees. We randomly select a fixed ratio of features to build each decision tree. Unless otherwise specified, we utilize half of the features for each weak learner and set $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. The run time is measured on a single Intel Xeon Gold 6226 2.9 Ghz CPU processor.
4.1 Cell-level outlier detection
Experimental setting
In practical situations, even when dealing with abnormal data points, it is not always the case that all cells are noisy [37, 29, 21]. To simulate more realistic settings, we introduce noise to certain cells in the following two-step process: First, we randomly select $20\%$ rows for each dataset. We then select $20\%$ columns uniformly at random, allowing each selected row to have a different set of perturbed cells. We inject noises sampled from the low-probability region into these cells, following Du et al. [8] and Liu et al. [29]. Details on the outlier injection process can be found in Appendix A.3.
We use $12$ publicly accessible binary classification datasets from OpenML, encompassing a range of both low and high-dimensional datasets, which have been widely used in the literature [12, 22, 23]. Details on these datasets are presented in Appendix A.1. For each dataset, $1000$ and $3000$ data points are randomly sampled for training and test datasets, respectively. For the baseline method, we consider 2D-KNN, a fast and performant variant of 2D-Shapley [29]. We incorporate a distance regularization term in the utility function $T$ for enhanced performance.
Results
We calculate the valuations for each cell using our joint valuation framework. Ideally, the outlier cells should receive a low valuation. We then arrange the cell valuations in ascending order and inspect those cells with the lowest values first.
The detection rate curve of inserted outlier is shown in Figure 2. For all datasets, 2D-OOB successfully identifies over $90\%$ of the outlier cells by inspecting only $30\%$ of the bottom cells. In comparison, 2D-KNN requires examining nearly $90\%$ of the cells to achieve the same detection level.
We also evaluate the area under the curve (AUC) as a quantitative metric and the run-time. As Table 1 shows, 2D-OOB achieves an average AUC of $0.83$ across $12$ datasets, compared to $0.67$ for 2D-KNN, while being significantly faster. For high-dimensional datasets such as the musk dataset, which comprises $166$ features, 2D-KNN would take more than an hour to process, while 2D-OOB can finish in seconds. Furthermore, we present additional results on multi-class classification datasets in Appendix B.1, demonstrating the consistently superior performance and efficiency of 2D-OOB.
Table 1: Cell-level outlier detection results. AUC and run-time comparison between 2D-OOB and 2D-KNN across the twelve datasets. The average and standard error of the AUC and run-time (in seconds) based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Bold numbers denote the best method. The AUC value for the Random method consistently remains at $0.5$ across all datasets. Overall, 2D-OOB achieves a significantly higher AUC while being orders of magnitude faster than 2D-KNN.
Dataset AUC $\uparrow$ Run-time $\downarrow$
2D-OOB (ours) 2D-KNN 2D-OOB (ours) 2D-KNN
lawschool 0.88$\pm$ 0.0027 0.75$\pm$ 0.0011 3.33 $\pm$ 0.06 177.56 $\pm$ 1.92
electricity 0.77$\pm$ 0.0072 0.68$\pm$ 0.0014 3.39 $\pm$ 0.07 191.38 $\pm$ 2.60
fried 0.91$\pm$ 0.0015 0.61$\pm$ 0.0005 3.97 $\pm$ 0.10 322.79 $\pm$ 2.98
2dplanes 0.87$\pm$ 0.0015 0.62$\pm$ 0.0005 3.46 $\pm$ 0.05 295.25 $\pm$ 2.37
creditcard 0.72$\pm$ 0.0028 0.69$\pm$ 0.0011 4.56 $\pm$ 0.10 662.34 $\pm$ 7.12
pol 0.82$\pm$ 0.0014 0.67$\pm$ 0.0006 4.34 $\pm$ 0.05 759.33 $\pm$ 4.37
MiniBooNE 0.77$\pm$ 0.0058 0.63$\pm$ 0.0019 7.46 $\pm$ 0.06 1507.83 $\pm$ 14.50
jannis 0.83$\pm$ 0.0042 0.55$\pm$ 0.0004 7.98 $\pm$ 0.07 1753.10 $\pm$ 12.35
nomao 0.79$\pm$ 0.0021 0.67$\pm$ 0.0009 7.69$\pm$ 0.11 2564.58 $\pm$ 23.11
vehicle_sensIT 0.81$\pm$ 0.0014 0.61$\pm$ 0.0005 9.87 $\pm$ 0.08 3113.65 $\pm$ 24.54
gas_drift 0.86$\pm$ 0.0010 0.84$\pm$ 0.0017 11.28$\pm$ 0.10 3878.31 $\pm$ 40.72
musk 0.88$\pm$ 0.0008 0.71$\pm$ 0.0006 14.09 $\pm$ 0.11 4415.45 $\pm$ 22.96
Average 0.83 0.67 6.78 1636.80
4.2 Cell fixation experiment
Experimental setting
A naive strategy to handle cell-level outliers is to eliminate data points that contain outliers. This method, however, risks substantial data loss, particularly when outliers are scattered and data points are costly to collect. We instead consider a cell fixation experiment, where we assume that the ground-truth annotations of outlier cells can be restored with external expert knowledge. At each step, we “fix" a certain number of cells by substituting them with their ground-truth annotations, prioritizing cells that have the lowest valuations. Then we fit a logistic model and evaluate the model’s performance with a test set of $3000$ samples. It is important to note that correcting normal cells has no effect, whereas fixing outlier cells is expected to enhance the model’s performance. We adopt the same datasets and implementations as in Section 4.1.
Figure 3: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN, and Random. We replace cells with their corresponding ground-truth annotations, starting with those cells assigned the lowest valuations. The results from $6$ datasets are presented, and additional results are provided in Appendix B.2. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB demonstrates a superior capability in accurately identifying and rectifying cell-level outliers.
Results
Figure 3 illustrates the anticipated trend in the performance of 2D-OOB, validating our method’s capability to accurately identify and prioritize the most impactful outliers for correction. As cells with the lowest valuations are progressively fixed, 2D-OOB demonstrates a consistent improvement in model accuracy. In contrast, when applying the same procedure with 2D-KNN, such notable performance enhancements are not observed.
Additionally, we investigate a scenario where ground-truth annotations remain unavailable. We adopt the setup from Liu et al. [29], where we replace the outlier cells with the average of other cells in the same feature column. 2D-OOB uniformly demonstrates significant superiority over its counterparts. Results are provided in Appendix B.2.
Figure 4: Backdoor trigger detection rate curve for 2D-OOB, 2D-KNN, and Random. The panels A. and B. correspond to Trojan square and BadNets square, respectively. We prioritize cells within each poisoned sample, ranking from highest to lowest based on their valuations. The detection rate curve shows the average detection rate across all poisoned samples, and error bars represent a $95\%$ confidence interval based on 15 independent runs. 2D-OOB demonstrates superior performance in detecting the cells implanted with triggers.
4.3 Backdoor trigger detection
A common strategy of data poisoning attacks involves inserting a predefined trigger (e.g., a specific pixel pattern in an image) into a few training data [13, 5, 28]. These malicious manipulations can be challenging to detect as they only infect targeted samples. Even when poisoned data are present, it could be difficult to discern the cause of attacks since manually reviewing the images is expensive and time-consuming. In this experiment, we introduce a novel joint valuation task: detecting backdoor triggers in data poisoning attacks. Distinct from random outliers investigated previously, such cell contamination is targeted and deliberate.
Figure 5: Qualitative examples for 2D-OOB in the backdoor trigger detection task. Each pair of images shows a poisoned image and its cell valuation. The color of the heatmap indicates importance: red cells are more important than blue cells. The first two pairs consider the case the class “bird” is relabeled as “cat”, and the latter two pairs consider the case the class “deer” is relabeled as “cat”. The heatmaps clearly show that higher cell valuations predominantly concentrate on the regions containing triggers, while areas featuring actual objects receive lower valuations. This pattern suggests that 2D-OOB effectively captures the triggers as the impactful features responsible for the misclassification of the poisoned samples.
We consider two popular backdoor attack algorithms: BadNets [13] and Trojan Attack [28]. The poisoned samples, relabeled as the adversarial target class, are mixed up with the clean data in the training process. As a result, the model is trained to incorrectly treat the trigger as a main feature of the poisoned samples. At the test time, those inputs containing the trigger will be misclassified to the target class. In this context, our goal is to effectively pinpoint the triggers by recognizing them as influential features through our joint valuation framework.
Experimental setting We select $5$ pairs of CIFAR-10 classes. For each pair, we designate one as the target attack class and the other as the source class. The training dataset comprises $1000$ images. For each attack, we contaminate $15\%$ of the training samples from the source class and relabel them to the target class. Two types of attack triggers are implemented: Trojan square and BadNets square [13, 35, 28]. These triggers are placed in the lower right corner of the original images to minimize occlusion. Details of these attacks are available in the Appendix A.4. In our experiment, the ratio of poisoned cells is approximately $1\%$. We sample 25% features to build each weak learner.
Results
We adopt the same detection scheme and baseline methods as in Section 4.1. Ideally, the poisoned cell should receive a high valuation based on the fact that such data point has been relabeled. We plot the detection rate curves of five datasets as shown in Figure 4. 2D-OOB significantly outperforms 2D-KNN in detecting both types of triggers. Overall, 2D-OOB achieves an average AUC of $0.95$ across all datasets and attack types, compared to $0.83$ for 2D-KNN.
Qualitative examples
Figure 5 displays the heatmaps for poisoned samples based on cell valuations of 2D-OOB. Areas with higher cell valuations (marked as dark red color) precisely indicate the trigger location in these samples, illustrating the effectiveness of our detection. More examples are included in the Appendix B.3.
4.4 Ablation study
We conduct ablation studies on the cell-level outlier detection task, as outlined in Section 4.1, to examine the impact of the selection and number of weak learners on 2D-OOB estimations.
Selection of weak learners
Although our study primarily employs decision trees as weak learners, it is important to note that 2D-OOB is model-agnostic, enabling the use of any class of machine learning models as weak learners. We compare efficacy of decision trees, logistic regression, a single-layer MLP with 64 dimensions, and a two-layer MLP with 64 and 32 dimensions.
Table 2 presents a comparison of detection AUC across 12 datasets, indicating that 2D-OOB is not model-free. The selection of weak learners slightly affects the valuation results, with more complex models generally yielding better performance. Nonetheless, all variations of 2D-OOB outperform 2D-KNN, highlighting the significant advantages of the 2D-OOB approach.
Table 2: Ablation study results of weak learner types. The average and standard error of the AUC based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Results of 2D-KNN are added for comparison. Different types of weak learners lead to variations in the valuation results, with more complex models generally showing better performance.
Dataset Decision Tree Logistic Regression MLP (single-layer) MLP (two-layer) 2D-KNN (Baseline)
lawschool 0.88 $\pm$ 0.0027 0.81 $\pm$ 0.0014 0.83 $\pm$ 0.0023 0.86 $\pm$ 0.0049 0.75 $\pm$ 0.0011
electricity 0.77 $\pm$ 0.0072 0.75 $\pm$ 0.0029 0.75 $\pm$ 0.0039 0.74 $\pm$ 0.0064 0.68 $\pm$ 0.0014
fried 0.91 $\pm$ 0.0015 0.82 $\pm$ 0.0023 0.85 $\pm$ 0.0020 0.88 $\pm$ 0.0027 0.61 $\pm$ 0.0005
2dplanes 0.87 $\pm$ 0.0015 0.82 $\pm$ 0.0026 0.86 $\pm$ 0.0026 0.88 $\pm$ 0.0037 0.62 $\pm$ 0.0005
creditcard 0.72 $\pm$ 0.0028 0.74 $\pm$ 0.0023 0.74 $\pm$ 0.0026 0.74 $\pm$ 0.0071 0.69 $\pm$ 0.0011
pol 0.82 $\pm$ 0.0014 0.79 $\pm$ 0.0029 0.85 $\pm$ 0.0014 0.86 $\pm$ 0.0019 0.67 $\pm$ 0.0006
MiniBooNE 0.77 $\pm$ 0.0058 0.77 $\pm$ 0.0059 0.80 $\pm$ 0.0057 0.81 $\pm$ 0.0119 0.63 $\pm$ 0.0019
jannis 0.83 $\pm$ 0.0042 0.76 $\pm$ 0.0040 0.79 $\pm$ 0.0048 0.80 $\pm$ 0.0108 0.55 $\pm$ 0.0004
nomao 0.79 $\pm$ 0.0021 0.82 $\pm$ 0.0012 0.83 $\pm$ 0.0010 0.83 $\pm$ 0.0017 0.67 $\pm$ 0.0009
vehicle-sensIT 0.81 $\pm$ 0.0014 0.81 $\pm$ 0.0026 0.80 $\pm$ 0.0025 0.82 $\pm$ 0.0037 0.61 $\pm$ 0.0005
gas-drift 0.86 $\pm$ 0.0010 0.89 $\pm$ 0.0005 0.88 $\pm$ 0.0005 0.88 $\pm$ 0.0006 0.84 $\pm$ 0.0017
musk 0.88 $\pm$ 0.0008 0.87 $\pm$ 0.0005 0.88 $\pm$ 0.0005 0.88 $\pm$ 0.0008 0.71 $\pm$ 0.0006
Average 0.83 0.80 0.82 0.83 0.67
The number of weak learners
Increasing the number of weak learners allows for a greater number of data-feature subset pairs to be explored, potentially leading to more accurate estimates. However, we empirically observe that beyond a certain threshold, adding extra weak learners does not substantially enhance performance, indicating convergence of the estimation in Appendix B.4. As a summary, we vary the number of weak learners $B\in\{500,1000,3000\}$ and compare the cell-level outlier detection performance. Typically, when the number of weak learners is 1000, i.e., $B=1000$, it is sufficient to achieve converged estimates across different datasets.
Lastly, we present additional ablation study results for other key hyperparameters in Appendix B.4. Apart from the experiments discussed above, we showcase that marginalization of 2D-OOB can either match or surpass state-of-the-art data valuation methods on standard benchmarks in Appendix D.
5 Related work
Data contribution estimation
In addition to the marginal contribution-based methods discussed in Section 2, many other approaches are emerging in the area of data valuation. Just et al. [18] develop a non-conventional class-wise Wasserstein distance between the training and validation sets and use the gradient information to evaluate each data point. Wu et al. [47] extend data valuation to deep neural networks, introducing a training-free data valuation framework based on neural tangent kernel theory. Yoon et al. [48] leverage reinforcement learning techniques to automatically learn data valuation scores by training a regression model. However, all these data valuation methods do not assign importance scores to cells, whereas our method provides additional insights into how individual cells contribute to the data valuations.
Feature attribution
Feature attribution is a pivotal research domain in explainable machine learning that primarily aims to provide insights into how individual features influence model predictions. Various effective methods have been proposed, including SHAP-based explanation [30, 31, 24, 7, 6], counterfactual explanation [44, 17, 36, 32, 33], and backpropagation-based explanation [1, 2, 42, 41, 49]. Among these methods, the SHAP-based explanation stands out as the most widely adopted approach, utilizing cooperative game theory principles to compute the Shapley value [38]. While feature attribution offers a potential method to attribute data valuation scores across individual cells, our empirical experiments in Appendix B.1 reveal that this two-stage scheme falls short in efficacy compared to our proposed joint valuation paradigm, which integrates data valuation and feature attribution in a simultaneous process.
6 Conclusion
We propose 2D-OOB, an efficient joint valuation framework that assigns a score to each cell in a dataset, thereby facilitating a finer attribution of data contribution and enabling a deeper understanding of datasets. Through comprehensive experiments, we show that 2D-OOB is computationally efficient and competitive over state-of-the-art methods in both joint valuation tasks.
Limitation and future work
While our study primarily explores random forest models applied to tabular datasets and simple image datasets, the potential application of neural network models within the 2D-OOB framework for more complex vision and language tasks presents a promising avenue for future investigation. For instance, in text datasets, tokens or words can be treated as cells. 2D-OOB can be easily integrated into any bagging training scheme that uses language models.
Overall, we believe that our work will inspire further exploration in the field of joint valuation, with the broader goal of improving the transparency and interpretability of machine learning, as well as developing an equitable incentive mechanism for data sharing.
Supplementary Materials
In the supplementary materials, we provide implementation details, additional experimental results, rigorous formalized proofs and data valuation experiment results. Code repository can be found at https://github.com/YifanSun99/2d-oob.
Appendix A Implementation details
A.1 Datasets
Tabular datasets
We use $12$ binary classification datasets obtained from OpenML [11]. A summary of all the datasets is provided in Table 3. These datasets are used in Section 4.14.2, and Appendix D.
For each dataset, we first employ a standard normalization procedure, where each feature is normalized to have zero mean and unit standard deviation. After preprocessing, we randomly partition a subset of the data into two non-overlapping sets: a training dataset and a test dataset, which consists of $1000$ and $3000$ samples respectively. The training dataset is used to obtain the joint (or marginal) valuation for each cell (or data point). The test dataset is exclusively used for cell fixation (or point removal) experiments when evaluating the test accuracy. Note that for methods that need a validation dataset such as KNNShapley and DataShapley, we additionally sample a separate validation dataset (disjoint from training dataset and test dataset) to evaluate the utility function. The size of the validation dataset is set to $10\%$ of the training sample size.
Image datasets
We create datasets by pairing CIFAR-10 classes, each pair consisting of a target attack class and a source class. The training and test dataset comprises $1000$ and $2000$ samples respectively. The size of the validation dataset is set to $10\%$ of the training sample size. To manage the computational challenges posed by the baseline method, we employ the super-pixel technique to transform the (32,32,3) image into a $256$-dimensional vector. Specifically, we first average the pixel values across three channels for each pixel. Then, we partition these transformed images into equally sized $2\times 2$ grids. In each grid, we use average pooling to reduce the pixel values to a single cell value. These cell values are then arranged into a flattened input vector. We annotate a cell as poisoned if at least $25\%$ of its corresponding grid area contains the trigger.
A.2 Implementation details for different methods
2D-OOB
2D-OOB involves fitting a subset random forest model with $B=1000$ decision trees based on the package “scikit-learn”. When constructing each decision tree, we fix the feature subset size ratio as $0.5$. Ablations on the hyperparameters can be found in Appendix B.4. For Section 4.3 and Appendix D, we simply adopt $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. For Section 4.1 and 4.2, we further calculate the normalized negative L2 distance between covariates and the class-specific mean in the bootstrap dataset, denoted as $d_{norm}$. Then we use $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))+d_{norm}$.
2D-KNN
2D-KNN employs KNN as a surrogate model to approximate 2D-Shapley. We set the number of nearest neighbors as $10$ and the number of permutations as $1000$. The hyperparameters are selected based on convergence behavior and we determine the run time until the values converge.
A.3 Implementation details for cell-level outlier generation
Following Du et al. [8] and Liu et al. [29], we replace a given cell with the outlier value. Here, the outlier value is randomly generated from the two-sided “tails” of the Gaussian distribution with the column mean and standard deviation, where the probability of the two-sided tail area is set to be $1$%. $4\%$ ($20\%\times 20\%$) of the cells in total are replaced with the corresponding outlier value.
A.4 Implementation details for backdoor trigger generation
Following the prior work [13, 28], we generate the BadNets square and the Trojan square trigger. For BadNets, we adopt the implementation in Nicolae et al. [34]. For Trojan Attack, we use a pretrained ResNet18 model on CIFAR-10 dataset and employ the implementation in Pang et al. [35]. For each attack, we evaluate its effectiveness by training a decision tree model on the poisoned dataset. The accuracy on a clean test set remains nearly unchanged compared to the model trained on an uncontaminated training set, while the attack success rate on a hold-out poisoned test sample set is guaranteed to exceed $75\%$.
Table 3: A summary of all the datasets used in 4.14.2, and Appendix  D. These datasets have been commonly used in previous literature [12, 22, 23]
Name Total sample size Input dimension Majority class proportion OpenML ID
lawschool 20800 6 0.679 43890
electricity 38474 6 0.5 44080
fried 40768 10 0.502 901
2dplanes 40768 10 0.501 727
creditcard 30000 23 0.779 42477
pol 15000 48 0.664 722
MiniBooNE 72998 50 0.5 43974
jannis 57580 54 0.5 43977
nomao 34465 89 0.715 1486
vehicle_sensIT 98528 100 0.50 357
gas_drift 5935 128 0.507 1476
musk 6598 166 0.846 1116
Appendix B Additional experimental results
In Section 4.1, we demonstrate that 2D-OOB shows promising performance in identifying cell-level outliers. This section further shows that our result is not sensitive to the selection of hyperparameters. Furthermore, ours generally performs better than 2D-KNN in different settings. We also provide additional results for Section 4.2 and  4.3.
B.1 Additional results for cell-level outlier detection
Additional results on multi-class classification datasets
We have conducted cell-level outlier detection experiments (as in Section 4.1) on three multi-class classification datasets from the UCI Machine Learning repository [19]. As shown in the table, 2D-OOB displays superior detection performance and efficiency.
Table 4: Cell-level outlier experiment results on multi-class classification datasets The average and standard error of the detection AUC and Elapsed Time (in seconds) based on 30 independent experiments are denoted by “average $\pm$ standard error“.
Dataset AUC $\uparrow$ Run-time $\downarrow$
2D-OOB (ours) 2D-KNN 2D-OOB (ours) 2D-KNN
Covertype 0.81$\pm$0.0156 0.63$\pm$0.0183 3.98$\pm$0.5774 962.34$\pm$1.3383
Dry Bean 0.88$\pm$0.0059 0.85$\pm$0.0192 3.31$\pm$0.4586 347.80$\pm$2.0212
Wine Quality 0.86$\pm$0.0178 0.57$\pm$0.0252 2.90$\pm$0.1240 269.14$\pm$1.1825
Additional baseline: two-stage attribution
Once we obtain the data valuation scores, an alternative solution approach to determining cell-level attributions involves leveraging feature attribution methods such as SHAP [30]. We explore an additional baseline method building upon this idea: initially, Data-OOB (or any other data valuation method) is computed for the $i$-th data point, denoted as $dv_{i}$. Subsequently, TreeSHAP [31] is fitted, using $dv_{i}$ as the target and the concatenation of $x_{i}$ and $y_{i}$ (denoted as $x_{i}\oplus y_{i}$) as the predictor. The derived local feature attributions are then interpreted as joint valuation results. We refer to this method as ‘two-stage attribution’.
Table 5 indicates that 2D-OOB substantially outperforms its two-stage counterpart. We hypothesize that the superiority of our method stems from integrating data valuation and feature attribution into a cohesive framework. Conversely, the two-stage method treats data valuation and feature attribution as separate processes, potentially resulting in sub-optimal outcomes. Furthermore, due to the computational complexity of TreeSHAP, the two-stage approach is notably slower compared to our method.
Table 5: Cell-level outlier detection results (AUC) of 2D-OOB and the two-stage attribution. Our method shows a better performance than the alternative method by a significant performance margin.
Dataset AUC $\uparrow$
2D-OOB (ours) Two-stage attribution
lawschool 0.88$\pm$ 0.0027 0.83$\pm$ 0.0064
electricity 0.77$\pm$ 0.0072 0.64$\pm$ 0.0093
fried 0.91$\pm$ 0.0015 0.82$\pm$ 0.0068
2dplanes 0.87$\pm$ 0.0015 0.80$\pm$ 0.0058
creditcard 0.72$\pm$ 0.0028 0.67$\pm$ 0.0051
pol 0.82$\pm$ 0.0014 0.78$\pm$ 0.0042
MiniBooNE 0.77$\pm$ 0.0058 0.70$\pm$ 0.0041
jannis 0.83$\pm$ 0.0042 0.62$\pm$ 0.0043
nomao 0.79$\pm$ 0.0021 0.71$\pm$ 0.0041
vehicle_sensIT 0.81$\pm$ 0.0014 0.64$\pm$ 0.0033
gas_drift 0.86$\pm$ 0.0010 0.73$\pm$ 0.0143
musk 0.88$\pm$ 0.0008 0.68$\pm$ 0.0028
Average 0.83 0.72
A noisy setting with more outlier cells
We consider a more challenging scenario with increased outlier levels, where both the row outlier ratio and column outlier ratio increase from $20\%$ (as in Section 4.1) to $50\%$. Consequently, this leads to $25\%$ ($50\%\times 50\%$) of the cells being replaced with outlier values. We follow the same outlier generation procedure outlined in Appendix A.3. The findings, presented in Table 6, demonstrate that our method maintains a significantly superior performance over 2D-KNN, even under such a noisy setting.
Table 6: Cell-level outlier detection results (AUC) of different joint valuation methods when the row outlier ratio and column outlier ratio are both $50\%$. Our method consistently outperforms 2D-KNN even in the presence of significant noise.
Dataset AUC $\uparrow$
2D-OOB (ours) 2D-KNN
lawschool 0.75$\pm$ 0.0084 0.60$\pm$ 0.0144
electricity 0.64$\pm$ 0.0155 0.60$\pm$ 0.0106
fried 0.74$\pm$ 0.0087 0.54$\pm$ 0.0027
2dplanes 0.74$\pm$ 0.0063 0.55$\pm$ 0.0033
creditcard 0.63$\pm$ 0.0055 0.61$\pm$ 0.0053
pol 0.69$\pm$ 0.0069 0.60$\pm$ 0.0042
MiniBooNE 0.67$\pm$ 0.0128 0.60$\pm$ 0.0048
jannis 0.70$\pm$ 0.0113 0.53$\pm$ 0.0014
nomao 0.70$\pm$ 0.0088 0.58$\pm$ 0.0052
vehicle_sensIT 0.70$\pm$ 0.0075 0.55$\pm$ 0.0031
gas_drift 0.73$\pm$ 0.0077 0.65$\pm$ 0.0114
musk 0.77$\pm$ 0.0063 0.64$\pm$ 0.0038
Average 0.71 0.59
B.2 Additional results for cell fixation experiment
Figure 6 presents the results for the cell fixation experiment on $6$ additional datasets. 2D-OOB excels in precisely detecting and correcting relevant cell outliers.
Figure 6: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with ground-truth values from the cells with the lowest valuation to the highest valuation. The results from $6$ datasets are displayed. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB sets itself apart by its remarkable precision in detecting and rectifying relevant cell outliers.
The scenario without ground-truth knowledge
Following [29], we examine a situation where external information on the ground-truth annotations of outlier cells is not accessible. In this scenario, we address these outliers by substituting them with the average of other cells in the same feature column. This procedure starts by addressing cells with the lowest valuations, based on the hypothesis that correcting these cells is likely to maintain or potentially improve the model’s performance. As depicted in Figure 7, 2D-OOB conforms to this expected trend, demonstrating the effectiveness of our method in joint valuation. Conversely, 2D-KNN fails to show similar performance improvements.
Figure 7: Cell fixation experiment (without ground-truth knowledge) results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with column mean imputations from the cells with the lowest value to the highest value. The results from $6$ datasets are displayed. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance.
B.3 Additional results for backdoor trigger experiment
We provide additional qualitative examples of backdoor trigger detection experiments in Figure 8.
B.4 Additional results for ablation study
We present the results of ablation study on the number of base learners $B$ and feature subset ratio $K/d$. Specifically, we examine the AUC of the detection curve in the cell-level outlier detection experiment (refer to Table 1).
The number of base learners $B$
When we increase the number of base learners from $500$ to $3000$, the detection AUC for each dataset remains unchanged, as shown in Table 7. This indicates that $1000$ base learners are sufficient to get an equitable joint valuation.
Table 7: Ablation results on the number of base learners $B$. The cell-level outlier detection results (AUC) are examined. Increasing the number of base learners from $1000$ to $3000$ does not yield a notable performance improvement.
Dataset AUC $\uparrow$
$B=500$ $B=1000$ $B=3000$
lawschool 0.86 $\pm$ 0.0035 0.88 $\pm$ 0.0027 0.88 $\pm$ 0.0026
electricity 0.77 $\pm$ 0.0062 0.77 $\pm$ 0.0072 0.77 $\pm$ 0.0070
fried 0.87 $\pm$ 0.0022 0.91 $\pm$ 0.0015 0.91 $\pm$ 0.0014
2dplanes 0.87 $\pm$ 0.0016 0.87 $\pm$ 0.0015 0.87 $\pm$ 0.0015
creditcard 0.72 $\pm$ 0.0025 0.72 $\pm$ 0.0028 0.72 $\pm$ 0.0028
pol 0.78 $\pm$ 0.0022 0.82 $\pm$ 0.0014 0.82 $\pm$ 0.0014
MiniBooNE 0.77 $\pm$ 0.0042 0.77 $\pm$ 0.0058 0.77 $\pm$ 0.0058
jannis 0.78 $\pm$ 0.0045 0.83 $\pm$ 0.0042 0.83 $\pm$ 0.0039
nomao 0.79 $\pm$ 0.0018 0.79 $\pm$ 0.0021 0.79 $\pm$ 0.0020
vehicle_sensIT 0.80 $\pm$ 0.0021 0.81 $\pm$ 0.0014 0.81 $\pm$ 0.0014
gas_drift 0.86 $\pm$ 0.0007 0.86 $\pm$ 0.0010 0.86 $\pm$ 0.0010
musk 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008
Average 0.81 0.83 0.83
Feature subset ratio $K/d$
In addition to $0.50$, We test two additional feature subset ratios $0.25$ and $0.75$. The results in Table 8 suggest that in general, the joint valuation capacity of our method is robust to the choice of feature subset ratio.
Table 8: Ablation results on feature subset ratio $K/d$. The cell-level outlier detection results (AUC) are examined. Our method’s joint valuation capacity remains relatively stable regardless of the selected feature subset ratio.
Dataset AUC $\uparrow$
$K/d=0.25$ $K/d=0.50$ $K/d=0.75$
lawschool 0.86$\pm$ 0.0026 0.88$\pm$ 0.0027 0.88$\pm$ 0.0024
electricity 0.79$\pm$ 0.0070 0.77$\pm$ 0.0072 0.73$\pm$ 0.0070
fried 0.86$\pm$ 0.0024 0.91$\pm$ 0.0015 0.89$\pm$ 0.0007
2dplanes 0.82$\pm$ 0.0015 0.87$\pm$ 0.0015 0.88$\pm$ 0.0014
creditcard 0.73$\pm$ 0.0029 0.72$\pm$ 0.0028 0.71$\pm$ 0.0028
pol 0.66$\pm$ 0.0031 0.82$\pm$ 0.0014 0.82$\pm$ 0.0014
MiniBooNE 0.78$\pm$ 0.0076 0.77$\pm$ 0.0058 0.77$\pm$ 0.0049
jannis 0.84$\pm$ 0.0035 0.83$\pm$ 0.0042 0.82$\pm$ 0.0043
nomao 0.79$\pm$ 0.0019 0.79$\pm$ 0.0021 0.78$\pm$ 0.0021
vehicle_sensIT 0.81$\pm$ 0.0014 0.81$\pm$ 0.0014 0.80$\pm$ 0.0015
gas_drift 0.88$\pm$ 0.0009 0.86$\pm$ 0.0010 0.86$\pm$ 0.0009
musk 0.89$\pm$ 0.0008 0.88$\pm$ 0.0008 0.88$\pm$ 0.0008
Average 0.81 0.83 0.82
Figure 8: Qualitative results on more datasets for the backdoor trigger detection experiment. The first two images originate from the class “airplane” while relabeled as “automobile”. The latter two images originate from the class “frog” while relabeled as “horse”.
Appendix C Proof of Proposition 3.1
Proof.
For simplicity, we denote $\phi^{\mathrm{OOB}}_{i}(S)$ as $\phi_{i}(S)$ and $\psi^{\mathrm{2D-OOB}}_{ij}$ as $\psi_{ij}$ in the proof. With the set of subsets $\mathcal{S}:=\{S\subseteq[d]\}$ and the definition Data-OOB $\phi_{i}(S)$ for all $i\in[n]$, where $S$ is a feature subset, we denote the cardinality of $\mathcal{S}$ as $L:=|\mathcal{S}|=2^{d}$. Let $\gamma_{b}$ be a weight vector $\gamma_{b}:=(\gamma_{b1},\ldots,\gamma_{bL})$ for all $b\in[B]$, where $\gamma_{bl}\in\{0,1\}$ and $\gamma_{bl}=1$ indicates the $l$-th subset is used in the $b$-th weak learner. With $\{w_{b},\gamma_{b},\hat{f}_{b}\}_{b=1}^{B}$, we can also denote the $i$-th Data-OOB on $l$-th feature subset $S_{l}$ as
$\phi_{i}(S_{l})=\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)T(y_{i},\hat{f}_{b}(x_{i,S_{l}}))}{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)}.$
With slight abuse of notation, the formulation of 2D-OOB in  (4) can be expressed as follows.
$\psi_{ij}$ $=\frac{\sum_{l=1}^{L}\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)\mathds{1}(j\in S_{l})T(y_{i},\hat{f}_{b}(x_{i,S_{l}}))}{\sum_{l=1}^{L}\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)\mathds{1}(j\in S_{l})}$
$=\sum_{l=1}^{L}\mathds{1}(j\in S_{l})\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)T(y_{i},\hat{f}_{b}(x_{i,S_{l}}))}{\sum_{l=1}^{L}\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)\mathds{1}(j\in S_{l})}$
$=\sum_{l=1}^{L}\mathds{1}(j\in S_{l})\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)}{\sum_{l=1}^{L}\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)\mathds{1}(j\in S_{l})}\frac{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)T(y_{i},\hat{f}_{b}(x_{i,S_{l}}))}{\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1)}$
$=\sum_{l=1}^{L}\alpha_{i,j,l}\phi_{i}(S_{l}),$
where $\alpha_{i,j,l}\propto\mathds{1}(j\in S_{l})\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1),\forall i\in[n],j\in[d],l\in[L]$ and $\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Define $P_{i}(S_{l}|j\in S_{l},\{w_{bi}\}_{b=1}^{B})=\alpha_{i,j,l}$, which specifies an empirical distribution of the feature subset $S$, conditioned on $j\in S$, in relation to the bootstrap sampling process. Here, $\mathds{1}(j\in S_{l})$ implies the distribution is conditioned on the presence of the $j$-th feature within the feature subset $S_{l}$. $w_{bi}$ indicates whether the $i$-th sample is out-of-bag in the $b$-th bootstrap, and $\gamma_{bl}$ indicates whether the $l$-th feature subset is selected in the $b$-th weak learner. Thus, the point mass is determined by the sampling process.
Appendix D Data valuation experiment
In this section, we show that 2D-OOB-data, the marginalization of 2D-OOB, offers an effective approach to data valuation. This serves as the basis of our enhanced performance in joint valuation.
Marginalization
2D-OOB aims to attribute data contribution through cells. Consequently, by summing up 2D-OOB over all columns, we can derive data contribution values. For $i\in[n]$, we define the 2D-OOB-data $\psi^{data}_{i}$ as follows.
$\psi^{data}_{i}:=\frac{1}{d}\sum_{j=1}^{d}\psi^{\mathrm{2D-OOB}}_{ij},$ (5)
Proof.
Based on definition of 2D-OOB-Data, for $i\in[n]$,
$\psi^{data}_{i}:=\frac{1}{d}\sum_{j=1}^{d}\psi^{\mathrm{2D-OOB}}_{ij}$ $=\frac{1}{d}\sum_{j=1}^{d}\sum_{l=1}^{L}\alpha_{i,j,l}\phi^{\mathrm{OOB}}_{i}(S_{l})$
$=\sum_{l=1}^{L}(\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l})\phi^{\mathrm{OOB}}_{i}(S_{l}),$
where $\alpha_{i,j,l}$ is defined in Appendix C. We have $\sum_{l=1}^{L}(\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l})=\frac{1}{d}\sum_{j=1}^{d}\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Denote $P_{i}(S_{l}|\{w_{bi}\}_{b=1}^{B})=\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l}$, which induces the empirical expectation of Data-OOB with respect to $S_{l}$. ∎
Based on discussions in Section 3.2, the marginalizations also connect with Data-OOB:
Proposition D.1.
For all $i\in[n]$, the marginalizations $\psi^{data}_{i}$ can be expressed as follows.
$\psi^{data}_{i}$ $=\mathbb{E}_{\hat{F}_{S}}[\phi^{OOB}_{i}(S)],$
where the notations follow the same definitions as Proposition 3.1.
Proposition D.1 indicates 2D-OOB-data $\psi^{data}_{i}$ can be expressed as the average Data-OOB value for the $i$-th data point. As a result, 2D-OOB-data is expected to inherit the advanced ability of Data-OOB in terms of data valuation, as will be empirically examined in the Appendix D.
Experimental setting
Following the standard protocol in Kwon and Zou [22, 23] and Jiang et al. [16], we randomly select $10$% of the data points and change its label to the other class. For joint valuation methods, we calculate the valuation of each cell and perform the marginalization over features to obtain the data valuation scores. For the baseline methods, we further incorporate several state-of-the-art data valuation methods including DataShapley [12], KNNShapley [15], DataBanzhaf [46], LAVA [18], and Data-OOB [23]. Implementation details are listed below. To guarantee a fair comparison, we also employ the decision tree as the base model in DataShapley and DataBanzhaf. Mislabeled data detection and data removal experiment are examined based on this setting. We adopt the same $12$ datasets as outlined in Section 4.1.
Data-OOB
Data-OOB involves fitting a random forest model without feature subset sampling, consisting of $1000$ decision trees.
DataShapley
We use a Monte Carlo-based algorithm. The Gelman-Rubin statistics is computed to determine the termination criteria of the algorithm. Following Jiang et al. [16], We adopt the threshold to be $1.05$. To ensure a fair comparison with the proposed method, we employ the decision tree model for the utility evaluation.
KNNShapley
We set the number of nearest neighbors to be $10\%$ of the sample size following Jia et al. [15].
LAVA
We calculate the class-wise Wasserstein distance following Just et al. [18]. The “OTDD” framework is adopted to complete the optimal transport calculation.
DataBanzhaf
We adopt the implementation from Jiang et al. [16]. We employ the decision tree model and set “the number of models to train” to $1000$.
D.1 Mislabeled data detection
We calculate the precision-recall curve by comparing the actual annotations, which denote whether data points are mislabeled, against the data valuation scores computed by different methods. Mislabeled data typically have a detrimental impact on model performance. Therefore, data points that receive a lower valuation score are regarded as having a higher chance of being mislabeled. We then determine AUCPR (the AUC of the precision-recall curve) as a quantitative metric to assess the detection efficacy.
As shown in Table 9, 2D-OOB-data consistently outperforms 2D-KNN-data across all datasets, suggesting its superior ability to detect mislabeled data points. It is worth noting that 2D-OOB-data’s results are on par with Data-OOB, while significantly exceeding the performance of other data valuation methods. These results are in line with our theoretical analysis regarding the resemblance between Data-OOB and 2D-OOB-data. However, it is important to highlight that applying Data-OOB to the joint tasks is not feasible as mentioned earlier, underscoring the necessity for the development of 2D-OOB.
Table 9: Point-level mislabeled data detection results. AUCPR of different data valuation and (marginalized) joint valuation methods. The average and standard error of the AUCPR based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Bold numbers denote the best method, for data valuation and joint valuation respectively. The AUCPR value for the Random method consistently remains at 0.5 across all datasets. 2D-OOB-data exhibits performance comparable to Data-OOB, while significantly surpassing 2D-KNN-data (the marginalization of 2D-KNN) and all other data valuation methods.
Dataset Data Valuation Joint Valuation (Marginalized)
KNNShapley LAVA DataBanzhaf DataShapley Data-OOB 2D-KNN-data 2D-OOB-data (ours)
lawschool 0.66 $\pm$ 0.013 0.13 $\pm$ 0.003 0.46 $\pm$ 0.008 0.88 $\pm$ 0.007 1.00 $\pm$ 0.000 0.46 $\pm$ 0.011 0.99 $\pm$ 0.002
electricity 0.22 $\pm$ 0.008 0.11 $\pm$ 0.002 0.18 $\pm$ 0.005 0.26 $\pm$ 0.007 0.44 $\pm$ 0.007 0.20 $\pm$ 0.006 0.39 $\pm$ 0.007
fried 0.40 $\pm$ 0.014 0.11 $\pm$ 0.002 0.22 $\pm$ 0.007 0.35 $\pm$ 0.009 0.76 $\pm$ 0.007 0.34 $\pm$ 0.010 0.73 $\pm$ 0.008
2dplanes 0.46 $\pm$ 0.016 0.12 $\pm$ 0.002 0.32 $\pm$ 0.007 0.54 $\pm$ 0.009 0.78 $\pm$ 0.008 0.44 $\pm$ 0.011 0.68 $\pm$ 0.010
creditcard 0.37 $\pm$ 0.007 0.11 $\pm$ 0.003 0.16 $\pm$ 0.004 0.28 $\pm$ 0.006 0.40 $\pm$ 0.007 0.20 $\pm$ 0.005 0.40 $\pm$ 0.007
pol 0.19 $\pm$ 0.017 0.11 $\pm$ 0.002 0.37 $\pm$ 0.010 0.58 $\pm$ 0.012 0.93 $\pm$ 0.004 0.29 $\pm$ 0.018 0.87 $\pm$ 0.005
MiniBooNE 0.41 $\pm$ 0.013 0.13 $\pm$ 0.006 0.23 $\pm$ 0.007 0.41 $\pm$ 0.010 0.78 $\pm$ 0.007 0.36 $\pm$ 0.008 0.78 $\pm$ 0.007
jannis 0.20 $\pm$ 0.007 0.11 $\pm$ 0.002 0.14 $\pm$ 0.003 0.17 $\pm$ 0.005 0.38 $\pm$ 0.010 0.19 $\pm$ 0.006 0.37 $\pm$ 0.010
nomao 0.61 $\pm$ 0.012 0.14 $\pm$ 0.003 0.33 $\pm$ 0.010 0.58 $\pm$ 0.009 0.87 $\pm$ 0.006 0.33 $\pm$ 0.011 0.88 $\pm$ 0.005
vehicle_sensIT 0.22 $\pm$ 0.009 0.11 $\pm$ 0.002 0.21 $\pm$ 0.007 0.33 $\pm$ 0.011 0.56 $\pm$ 0.010 0.14 $\pm$ 0.005 0.56 $\pm$ 0.010
gas_drift 0.87 $\pm$ 0.013 0.16 $\pm$ 0.006 0.42 $\pm$ 0.009 0.75 $\pm$ 0.008 0.98 $\pm$ 0.002 0.88 $\pm$ 0.006 0.98 $\pm$ 0.002
musk 0.33 $\pm$ 0.010 0.11 $\pm$ 0.003 0.31 $\pm$ 0.007 0.47 $\pm$ 0.012 0.85 $\pm$ 0.005 0.21 $\pm$ 0.008 0.85 $\pm$ 0.005
Average 0.41 0.12 0.28 0.47 0.73 0.34 0.71
D.2 Point removal experiment
Removing low-quality data points has the potential to enhance model performance. Based on this idea, we employ the point removal experiment, a widely used benchmark in data valuation [23, 12, 22]. According to the calculated data valuation scores, we progressively remove data points from the dataset in ascending order. Specifically, we begin by removing the data points with the lowest data valuations. Each time we remove a datum, we fit a logistic model and use the held-out test set consisting of $3000$ instances to evaluate the model performance. The expected behavior is that the model performance will improve initially as the detrimental data points are gradually eliminated from the training process. Removing an excessive number of data points may result in a drastically altered dataset. Consequently, we opt to remove the bottom $20\%$ data points.
Test accuracy curves throughout the data removal process are shown for $12$ datasets (Figure 9). A higher curve signifies better performance in terms of data valuation. Overall, 2D-OOB-data demonstrates similar performance to Data-OOB, while significantly outperforming all other data valuation methods and the random baseline. When a few data points with poor quality are removed, the test performance of 2D-OOB-data exhibits an evident increase. However, such a positive trend does not apply to other popular data valuation methods including DataShapley and LAVA. These findings highlight the potential of 2D-OOB-data in selecting a subset of critical data points that can maintain model performance when the dataset is pruned.
Figure 9: Point removal experiment results (test accuracy curves) of $7$ data valuation methods – 2D-OOB-data, 2D-KNN-data, Data-OOB, LAVA, DataBanzhaf, DataShapley, KNNShapley and a random baseline. We remove data points from the lowest valuation to the highest valuation. The results from $6$ binary classification datasets are displayed. For each dataset, we conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB-data demonstrates superior ability in finding a set of helpful data points.